A recent report indicated that 67% of teens 12 to 19 play a video game online at least once a week. How large of a sample would a researcher need to estimate the true proportion of teens 12 to 19 who play a video game online at least once a week within 3% with a 95% confidence level?

1.) 844
2.) 4489
3.) 944
4.) 4144

Answer by Boreal(15235)

The SE is z(0.975)sqrt{p*(1-p)/sqrt(n)} and that is equal to 0.03
1.96*sqrt (0.67*0.33)/n=0.03
1.96 (sqrt(.2211))=0.03 sqrt (n)
1.96*0.4702=0.03 sqrt (n)
multiply on left and square both sides
0.8494=0.0009 *n
n=0.8494/0.0009=943.75 or 944.
Choice 3